Warner likes waffles for breakfast. How much energy is used by a waffle maker that has a power rating of .7501 kW and is operated for 3.6h ?
1
Expert's answer
2019-12-16T11:05:08-0500

E=P×t=(0.7501)×(3.6×3600)=97212.96 kJ